# On the fixed locus of the antisymplectic involution of an EPW cube

### v1

@@ 3. EPW cubes
Let $A\subseteq\textstyle{\bigwedge\hskip-2.56073pt^{3}}\hskip 0.56905pt{V}_{6}$ be a Lagrangian subspace with no decomposable vectors and such that the degeneracy locus ${\mathsf{Z}}_{A}^{\geq 4}$ is empty. The double cover $g_{A}\colon\widetilde{\mathsf{Z}}_{A}\to{\mathsf{Z}}_{A}$ defined in Equation (1) is then branched over the smooth threefold ${\mathsf{Z}}_{A}^{\geq 3}\subseteq{\mathsf{Z}}_{A}$ and the EPW cube $\widetilde{\mathsf{Z}}_{A}$ is a smooth hyper-Kähler manifold of K3[3]-type. The ample class $h\coloneqq g_{A}^{*}\mathscr{O}_{\operatorname{\mathsf{Gr}}(3,V_{6})}(1)$ satisfies $q(h)=4$.
The smooth threefold ${\mathsf{W}}_{A}\coloneqq g_{A}^{-1}({\mathsf{Z}}_{A}^{\geq 3})\subseteq\widetilde{\mathsf{Z}}_{A}$ is the fixed locus of the involution $\iota_{A}$ associated to $g_{A}$. In this section, using results and numerical relations proved in Section 2, we compute the Euler characteristic of ${\mathsf{W}}_{A}$ as well as the cohomology class $[{\mathsf{W}}_{A}]\in H^{6}({\mathsf{Z}}_{A},{\bf Q})$. Additionally, we determine the Chern classes of ${\mathsf{W}}_{A}$.
@@ 3.1. Euler characteristic of ${\mathsf{W}}_{A}$
where $+$ means the $\iota_{A}$-invariant part. We have
@@
[…]
@@
The fixed subspace for the action of $\iota_{A}^{*}$ on $H^{6}(X,{\bf Q})\cap V_{(3)}$ is equal to $(\operatorname{Sym}^{2}\!\bar{T}\cdot h)\oplus{\bf Q}h^{3}$, since $\iota^{*}_{A}$ acts on $\bar{T}$ as $-\operatorname{Id}$ and fixes $h$. Since $\bar{T}$ has dimension $22$, this fixed subspace has dimension 254.
We consider now the action of $\iota^{*}_{A}$ on $H^{6}(X,{\bf Q})\cap V_{(1,1)}$. By Lemma 2.2, this action is equal to ±$\pm\gamma_{(1,1)}$, where $\gamma_{(1,1)}$ is the action induced by the involution $\iota_{A}^{*}\oplus\operatorname{Id}$ on $V$. The fixed subspace for $\gamma_{(1,1)}$ is ${\bigwedge\hskip-2.56073pt^{2}}\hskip 0.56905pt{\bar{T}}\oplus\textstyle{\bigwedge\hskip-2.56073pt^{2}}\hskip 0.56905pt{U}$, of dimension 232, and the fixed space for $-\gamma_{(1,1)}$ is equal to $h\wedge\bar{T}$, of dimension 22.
@@
3.1.1. The class $[{\mathsf{W}}_{A}]$ in $H^{6}(\widetilde{Z}_{A},{\bf Z})$
The class $[{\mathsf{W}}_{A}]$ is a 3-dimensional Hodge class of $\widetilde{\mathsf{Z}}_{A}$ hence, by Equation (2.3), it can be written as
$[{\mathsf{W}}_{A}]=ah^{3}+bhc_{2}+c\eta,$
@@ 3.1.1. The class $[{\mathsf{W}}_{A}]$ in $H^{6}(\widetilde{Z}_{A},{\bf Z})$
$[{\mathsf{W}}_{A}]=\frac{5}{8}\left(3h^{3}-hc_{2}\right)\in H^{6}(\widetilde{\mathsf{Z}}_{A},{\bf Z}).$
@@
Since ${\mathsf{Z}}^{\geq 2}_{A}\cdot h^{3}=720$ ([IKKR, Section 2.3]), it follows from Proposition 2.5 that the class $[{\mathsf{W}}_{A}]\in H^{6}(\widetilde{\mathsf{Z}}^{\geq 2}_{A},{\bf Q})$ is $[{\mathsf{W}}_{A}]=\frac{15}{8}h^{3}-\frac{5}{8}hc_{2}+c\eta.$ Since ${\mathsf{W}}_{A}$ is Lagrangian, it satisfies
$-\chi_{\textnormal{top}}({\mathsf{W}}_{A})=[{\mathsf{W}}_{A}]^{2}=1200+4c^{2}.$
We compare this with the two possible values of $\chi_{\textnormal{top}}({\mathsf{W}}_{A})$ computed in Equation (12) to obtain that $4c^{2}$ is either 00 or $336$, hence $c=0$ and $\chi_{\textnormal{top}}({\mathsf{W}}_{A})=-1200$. ∎
@@
has rank 1 and kernel $h^{\perp}$ (it maps the symplectic form to 00 and since the class of the latter describes a dense open subset of the intersection of $h^{\perp}$ and a smooth quadric, it vanishes on $h^{\perp}$; it is nonzero since the restriction of an ample class (such as $h$) is ample). This implies that $[{\mathsf{W}}_{A}]$ is an atomic Lagrangian [Be, Theorem 1.8]. Notice that its class belongs to the Verbitsky component, according to what is expected in [Be, Proposition 3.5].
@@ 3.2. Chern classes
By the discussion after [DK, Theorem 5.6], and [DK, Lemma 4.3] (applied with $S=\operatorname{\mathsf{Gr}}(3,V_{6})$, $k=3$, $c_{1}(\mathscr{L})=c_{1}(\mathscr{A}_{1})=0$, and $c_{1}(\mathscr{A}_{2})=\mathscr{O}_{\operatorname{\mathsf{Gr}}(3,V_{6})}(-4)$, and giving an equality of Cartier divisor classes), we have $\omega_{{\mathsf{Z}}_{A}^{\geq 3}}=\mathscr{O}_{{\mathsf{Z}}_{A}^{\geq 3}}(2)$, so that ${\mathsf{Z}}_{A}^{\geq 3}$ (hence also $W_{A}$) is a threefold of general type.
@@
$\chi({\mathsf{W}}_{A},\mathscr{O}_{{\mathsf{W}}_{A}})=\frac{1}{24}\,c_{1}(T_{{\mathsf{W}}_{A}})c_{2}(T_{{\mathsf{W}}_{A}})=-\frac{1}{24}\,(4h^{3}+hc_{2})\cdot[W_{A}]=-130.$
@@ 4. A singular degeneration of $\widetilde{\mathsf{Z}}^{\geq 2}_{A}$
In this section, we construct singular degenerations of $\widetilde{\mathsf{Z}}_{A}$. Our aim is to study the fixed locus of $\iota_{A}$ following the degeneration argument used in [FMOS1] and [FMOS2]. One family of singular EPW cubes was studied in [Ri]: their periods dominate the Heegner divisor ${}^{[3]}\mathscr{D}_{4,12}^{(2)}$ in the period domain. For the family that we construct below, the periods dominate another Heegner divisor, ${}^{[3]}\mathscr{D}_{4,2}^{(2)}$, in the same period domain (see [Ri, Appendix A]) for the notation).
@@
By [BC, Proposition 4.2 and Theorem 1.1], the closed movable and nef cones of $S^{[3]}$ coincide and, by [D, Example 3.20], they are spanned by the classes $L$ and $2L-\delta$. The class $2L-\delta$, of square $4$ and divisibility $2$, is therefore nef and big, but nonample on $S^{[3]}$. Therefore, the extremal ray ${\bf R}_{>0}(2L-\delta)$ induces a projective contraction
@@
Let $\bar{L}$ be the ample line bundle on the (singular) variety $X$ such that $\pi^{*}\bar{L}=L$. The biregular antisymplectic involution $j^{[3]}$ of $S^{[3]}$ induced by the involution $j$ of $S$ acts trivially on $\operatorname{NS}(S^{[3]})={\bf Z}L\oplus{\bf Z}\delta$ (and by $-\operatorname{Id}$ on its orthogonal complement in $H^{2}(S^{[3]},{\bf Z})$), so it descends to an involution $\iota$ on $X$.
The data $S^{[3]}$, $L$, $\Delta$, and $j^{[3]}$ satisfy requirements (a)–(d) of [FMOS1, Section 2]. Hence, we can apply [FMOS1, Proposition 2.1 and Corollary 2.2] (as in [FMOS2, Section 3.1]) to obtain a family
@@
  • the fiber over 00 is the variety $X$ with the ample line bundle $\bar{L}$ and involution $\iota$;
  • for any $t\in D\smallsetminus\{0\}$, $(\mathscr{X}_{t},\mathscr{L}_{t})$ is a smooth EPW cube $\widetilde{\mathsf{Z}}_{A_{t}}$ (for some Lagrangian $A_{t}\subseteq\textstyle{\bigwedge\hskip-2.56073pt^{3}}\hskip 0.56905pt{V}_{6}$) and $\iota_{t}$ is the natural EPW involution $\iota_{A_{t}}$.
@@ 4.1. The special fiber of $\mathscr{X}\to D$
such that $L=\Theta(0,-L,0)$ and $\delta=\Theta(-1,0,-2)$ ([BM1, Section 13]).
In [Br], Bridgeland showed that $S^{[3]}$ is also isomorphic to the moduli space $M_{\sigma}(v)$ of $\sigma$-stable objects in $D^{b}(S)$, where $\sigma$ is a stability condition contained in the distinguished connected component $\operatorname{Stab}^{\dagger}(S)$ (see [B, Section 3] for the definition).
Contractions of moduli spaces of sheaves on K3 surfaces are described in [BM1]. Their main result is that the wall and chamber decomposition of the movable cone of $M_{\sigma}(v)$ corresponds to the wall and chamber decomposition induced by $v$ on $\operatorname{Stab}^{\dagger}(S)$.
Moreover, the contraction associated to a certain wall $\mathscr{W}\subseteq\operatorname{Stab}^{\dagger}(S)$ is characterized via [BM1, Theorem 5.7] in terms of the associated hyperbolic lattice $H_{\mathscr{W}}\in H^{*}_{\operatorname{alg}}(S,Z)$ (see [BM1, Proposition 5.1] for the definition of $H_{\mathscr{W}}$). By definition, the hyperbolic lattice $H_{\mathscr{W}}$ contains $v$, and $\mathscr{W}$ corresponds to the wall $\Theta(H_{\mathscr{W}}^{\perp}\cap v^{\perp})$ of the movable cone of $M_{\sigma}(v)$.
In particular, in our case, contractions of $S^{[3]}$ are given by the two extremal rays ${\bf R}_{\geq 0}L$ and ${\bf R}_{\geq 0}(2L-\delta)$ of $\operatorname{Nef}(S)$. The extremal ray ${\bf R}_{\geq 0}L$ is orthogonal to the isotropic class $w\coloneqq(0,0,-1)$, which satisfies $w\cdot v=1$. The induced contraction is the Hilbert–Chow morphism $S^{[3]}\to S^{(3)}$ ([BM1, Theorem 5.7(a)]).
@@
$\pi|_{\Delta}\colon\Delta\longrightarrow W,$
where the smooth hyper-Kähler fourfold $W$ is the moduli space of stable sheaves on $S$ with Mukai vector $(0,L,-4)$.
@@
We use the notation of [BM1]. Under the isomorphism $\Theta$, the class $2L-\delta$ is the image of the vector $w=(1,-2L,2)$. In particular, the hyperbolic lattice associated to the contraction induced by $2L-\delta$ is $H=\langle v,(1,-L,2)\rangle=w^{\perp}\subseteq H^{2}_{\operatorname{alg}}(S,{\bf Z})$. It is a lattice with intersection form $\begin{pmatrix}4&0\\ 0&-2\end{pmatrix}$, and contains the spherical class $s=(1,-L,2)$.
@@
Note that, by [Br, Lemma 6.2], every $(\alpha,\beta)\in{\bf R}_{>0}\times{\bf R}\smallsetminus\{(1,-1\}$ that satisfies Equation (16) defines a stability condition, hence the wall $\mathscr{W}$ is equal to
@@
We analyze the wall $\mathscr{W}$ using [BM1, Theorem 5.7]. We first remark that $H$ contains no isotropic classes: indeed a vector $u=xv+ys\in H$, with $x,y\in{\bf Z}$, has square $4x^{2}-2y^{2}$ which is never 00.
Moreover, the wall $\mathscr{W}$ is not totally semistable. Since $\mathscr{W}$ is nonisotropic, to check this, we only need to show that there are no effective spherical classes $u$ with $(u,v)<0$. A spherical class $u\in H$ is a vector $u=xv+ys$ with $x,y\in{\bf Z}$ such that $12x^{2}-y^{2}=-$ 1, and the condition $(u,v)<0$ is equal to $x<0$. By [BM1, Proposition 5.5], a class $u$ is effective if and only if $\mathop{\rm Re}\nolimits\left(\frac{Z_{\alpha,\beta}(u)}{Z_{\alpha,\beta}(v)}\right)>0$ for one (hence for all) $(\alpha,\beta)\in\mathscr{W}$. In particular, we can choose $\beta=-2$ and $\alpha=\sqrt{2}$.
Using [B, Equation (3)], we compute $Z_{\alpha,\beta}(v)=-2\beta(\beta+2+i\alpha)$ and $Z_{\alpha,\beta}(s)=-2(\beta+1)(\beta+2+i\alpha)$ for $(\alpha,\beta)\in\mathscr{W}$. We obtain
@@
Note that for each pair $(x,y)$ with $x<0$ that satisfies $12x^{2}-y^{2}=-$ 1, we have $x+\frac{y}{2}<0$, thus there are no effective spherical classes $u$ with $(u,v)<0$.
Therefore the wall $\mathscr{W}$ is nonisotropic not totally semistable, and the lattice $H$ contains the spherical class $s$, which is effective on $\mathscr{W}$ since $\frac{Z_{\sqrt{2},-2}(s)}{Z_{\sqrt{2},-2}(v)}=1/2$. By [BM1, Theorem 5.7(a)], the associated contraction $\pi$ is divisorial of Brill–Noether type. In particular, by [BM1, Lemma 7.1], $\pi$ contracts a divisor $\Delta$ of class $\Theta(s)=L-\delta$. Moreover, since $s$ is the Mukai vector of the line bundle $\mathscr{O}_{S}(-L)$, the divisor $\Delta$ is given by the condition $\operatorname{Hom}(\mathscr{O}_{S}(-L),\bullet)\neq 0$. In particular, this corresponds to length-3 subschemes $Z\subseteq S$ that are contained in some curve $C\in|L|$.
More precisely, following the proof of [BM1, Lemma 7.4], the image $W\coloneqq\pi(\Delta)$ is isomorphic to the moduli space $M_{\sigma_{0}}(a)$, where $a\coloneqq v-s=(0,L,-4)$ and $\sigma_{0}=\sigma_{\alpha_{0},\beta_{0}}$ is a stability condition lying on the wall $\mathscr{W}$. By [BM, Example 9.7], we have equality $M_{\sigma_{0}}(a)=M_{S}(a)$ (by choosing $A=1$, $B=\beta_{0}$), hence $W$ is a smooth hyper-Kähler fourfold that parametrizes sheaves of the form $i_{C,*}(\xi)$, where $i_{C}\colon C\hookrightarrow S$ is the inclusion of a curve $C\in|L|$ in $S$ and $\xi$ is a torsion-free rank $1$ sheaf on $C$ of degree $-3$.
Finally, the fiber of $\pi$ over a point $[A]\in M_{S}(a)$ is isomorphic to ${\bf P}(\operatorname{Ext}^{1}(\mathscr{O}_{S}(-L),A))$. Since
@@
If $C\in|L|$ is smooth (of genus 2), a torsion-free rank $1$ sheaf on $C$ of degree $-3$ is of the form $\mathscr{O}_{C}(-Z)$, with $Z\in C^{(3)}$. If $A\coloneqq i_{C,*}(\mathscr{O}_{C}(-Z))$, the preimage $\pi^{-1}([A])$ is equal to $|\mathscr{O}_{C}(Z)|$ (they have the same dimension by Riemann–Roch, and any degree $3$ effective divisor on $C$ equivalent to $Z$ defines a point in $S^{[3]}$ that is sent to $[A]$ via $\pi$). ∎
We keep the notation of the proof. Let $\sigma_{0}$ be a stability condition on the wall $\mathscr{W}$ defined in Equation (17). The moduli space $M_{\sigma_{0}}(v)$ has been constructed as a good moduli space in the sense of Alper (see [T, Theorem 3.3]). By [BM, Proposition 8.1], there exists a morphism

### v2

@@ 3. EPW cubes
Let $A\subseteq\textstyle{\bigwedge\thinspace^{3}}\thinspace{V}_{6}$ be a Lagrangian subspace with no decomposable vectors and such that the degeneracy locus ${\mathsf{Z}}_{A}^{\geq 4}$ is empty. The double cover $g_{A}\colon\widetilde{\mathsf{Z}}_{A}\to{\mathsf{Z}}_{A}$ defined in Equation (1) is then branched over the smooth threefold ${\mathsf{Z}}_{A}^{\geq 3}\subseteq{\mathsf{Z}}_{A}$ and the EPW cube $\widetilde{\mathsf{Z}}_{A}$ is a smooth hyper-Kähler manifold of K3[3]-type. The ample class $h\coloneqq g_{A}^{*}\mathscr{O}_{\operatorname{\mathsf{Gr}}(3,V_{6})}(1)$ satisfies $q(h)=4$.
The smooth threefold ${\mathsf{W}}_{A}\coloneqq g_{A}^{-1}({\mathsf{Z}}_{A}^{\geq 3})\subseteq\widetilde{\mathsf{Z}}_{A}$ is the fixed locus of the involution $\iota_{A}$ associated to $g_{A}$. In this section, using results and numerical relations proved in Section 2, we compute the Euler characteristic of ${\mathsf{W}}_{A}$ as well as the cohomology class ${\mathsf{W}}_{A}]\in H^{6}({\mathsf{Z}}_{A},{\bf Q})$. Additionally, we determine the Chern classes of ${\mathsf{W}}_{A}$.
@@ 3.1. Euler characteristic of ${\mathsf{W}}_{A}$
where ++ means the $\iota_{A}$-invariant part. We have
@@
$H^{6}(X,{\bf Q})\cap V_{(1,1)}=\textstyle{\bigwedge\thinspace^{2}}\thinspace{\bar{T}}\oplus(h\wedge\bar{T})\oplus\textstyle{\bigwedge\thinspace^{2}}\thinspace{U},$
@@
The fixed subspace for the action of $\iota_{A}^{*}$ on $H^{6}(X,{\bf Q})\cap V_{(3)}$ is equal to $(\operatorname{Sym}^{2}\!\bar{T}\cdot h)\oplus{\bf Q}h^{3}$, since $\iota^{*}_{A}$ acts on $\bar{T}$ as $\operatorname{Id}$ and fixes $h$. Since $\bar{T}$ has dimension $22$, this fixed subspace has dimension 254.
We consider now the action of $\iota^{*}_{A}$ on $H^{6}(X,{\bf Q})\cap V_{(1,1)}$. By Lemma 2.2, this action is equal to $\pm\gamma_{(1,1)}$, where $\gamma_{(1,1)}$ is the action induced by the involution $\iota_{A}^{*}\oplus\operatorname{Id}$ on $V$. The fixed subspace for $\gamma_{(1,1)}$ is ${\bigwedge\thinspace^{2}}\thinspace{\bar{T}}\oplus\textstyle{\bigwedge\thinspace^{2}}\thinspace{U}$, of dimension 232, and the fixed space for $\gamma_{(1,1)}$ is equal to $h\wedge\bar{T}$, of dimension 22.
@@
3.1.1. The class ${\mathsf{W}}_{A}]$ in $H^{6}(\widetilde{Z}_{A},{\bf Z})$
The class ${\mathsf{W}}_{A}]$ is a 3-dimensional Hodge class of $\widetilde{\mathsf{Z}}_{A}$ hence, by Equation (2.3), it can be written as
${\mathsf{W}}_{A}]=ah^{3}+bhc_{2}+c\eta,$
@@ 3.1.1. The class ${\mathsf{W}}_{A}]$ in $H^{6}(\widetilde{Z}_{A},{\bf Z})$
${\mathsf{W}}_{A}]=\frac{5}{8}\left(3h^{3}-hc_{2}\right)\in H^{6}(\widetilde{\mathsf{Z}}_{A},{\bf Z}).$
@@
Since ${\mathsf{Z}}^{\geq 2}_{A}\cdot h^{3}=720$ ([IKKR19, Section 2.3]), it follows from Proposition 2.6 that the class of ${\mathsf{W}}_{A}$ in $H^{6}(\widetilde{\mathsf{Z}}^{\geq 2}_{A},{\bf Q})$ is ${\mathsf{W}}_{A}]=\frac{15}{8}h^{3}-\frac{5}{8}hc_{2}+c\eta.$ Since ${\mathsf{W}}_{A}$ is Lagrangian, it satisfies
$\chi_{\textnormal{top}}({\mathsf{W}}_{A})=[{\mathsf{W}}_{A}]^{2}=1200+4c^{2}.$
We compare this with the two possible values of $\chi_{\textnormal{top}}({\mathsf{W}}_{A})$ computed in Equation (14) to obtain that $4c^{2}$ is either 0 or $336$, hence $c=0$ and $\chi_{\textnormal{top}}({\mathsf{W}}_{A})=-1200$. ∎
@@
has rank 1 and kernel $h^{\perp}$ (it maps the symplectic form to 0 and since the class of the latter describes a dense open subset of the intersection of $h^{\perp}$ and a smooth quadric, it vanishes on $h^{\perp}$; it is nonzero since the restriction of an ample class (such as $h$) is ample). This implies that ${\mathsf{W}}_{A}]$ is an atomic Lagrangian [Bec25, Theorem 1.8]. Notice that its class belongs to the Verbitsky component, according to what is expected in [Bec25, Proposition 3.5].
@@ 3.2. Chern classes
By the discussion after [DK20, Theorem 5.6], and [DK20, Lemma 4.3] (applied with $S=\operatorname{\mathsf{Gr}}(3,V_{6})$, $k=3$, $c_{1}(\mathscr{L})=c_{1}(\mathscr{A}_{1})=0$, and $c_{1}(\mathscr{A}_{2})=\mathscr{O}_{\operatorname{\mathsf{Gr}}(3,V_{6})}(-4)$, and giving an equality of Cartier divisor classes), we have $\omega_{{\mathsf{Z}}_{A}^{\geq 3}}=\mathscr{O}_{{\mathsf{Z}}_{A}^{\geq 3}}(2)$, so that ${\mathsf{Z}}_{A}^{\geq 3}$ (hence also ${\mathsf{W}}_{A}$) is a threefold of general type.
@@
$\chi({\mathsf{W}}_{A},\mathscr{O}_{{\mathsf{W}}_{A}})=\frac{1}{24}\,c_{1}(T_{{\mathsf{W}}_{A}})c_{2}(T_{{\mathsf{W}}_{A}})=-\frac{1}{24}\,(4h^{3}+hc_{2})\cdot[{\mathsf{W}}_{A}]=-130.$
@@ 4. A singular degeneration of $\widetilde{\mathsf{Z}}^{\geq 2}_{A}$
In this section, we construct singular degenerations of $\widetilde{\mathsf{Z}}_{A}$. Our aim is to study the fixed locus of $\iota_{A}$ following the degeneration argument used in [FMOS22] and [FMOS23]. One family of singular EPW cubes was studied in [Riz24]: their periods dominate the Heegner divisor $^{[3]}\mathscr{D}_{4,12}^{(2)}$ in the period domain. For the family that we construct below, the periods dominate another Heegner divisor, $^{[3]}\mathscr{D}_{4,2}^{(2)}$, in the same period domain (see [Riz24, Appendix A]) for the notation).
@@
By [BC22, Proposition 4.2 and Theorem 1.1], the closed movable and nef cones of $S^{[3]}$ coincide and, by [BM14a, Proposition 13.1], they are spanned by the classes $L$ and $2L-\delta$. The class $2L-\delta$, of square $4$ and divisibility $2$, is therefore nef and big, but nonample on $S^{[3]}$. Therefore, the extremal ray ${\bf R}_{>0}(2L-\delta)$ induces a projective contraction
@@
Let $\bar{L}$ be the ample line bundle on the (singular) variety $X$ such that $\pi^{*}\bar{L}=L$. The biregular antisymplectic involution $j^{[3]}$ of $S^{[3]}$ induced by the involution $j$ of $S$ acts trivially on $\operatorname{NS}(S^{[3]})={\bf Z}L\oplus{\bf Z}\delta$ (and by $\operatorname{Id}$ on its orthogonal complement in $H^{2}(S^{[3]},{\bf Z})$), so it descends to an involution $\iota$ on $X$.
The data $S^{[3]}$, $L$, $\Delta$, and $j^{[3]}$ satisfy requirements (a)–(d) of [FMOS22, Section 2]. Hence, we can apply [FMOS22, Proposition 2.1 and Corollary 2.2] (as in [FMOS23, Section 3.1]) to obtain a family
@@
  • the fiber over 0 is the variety $X$ with the ample line bundle $\bar{L}$ and involution $\iota$;
  • for any $t\in D\smallsetminus\{0\}$, $(\mathscr{X}_{t},\mathscr{L}_{t})$ is a smooth EPW cube $\widetilde{\mathsf{Z}}_{A_{t}}$ (for some Lagrangian $A_{t}\subseteq\textstyle{\bigwedge\thinspace^{3}}\thinspace{V}_{6}$) and $\iota_{t}$ is the natural EPW involution $\iota_{A_{t}}$.
@@ 4.1. The special fiber of $\mathscr{X}\to D$
such that $L=\Theta(0,-L,0)$ and $\delta=\Theta(-1,0,-2)$ ([BM14a, Section 13]).
In [Bri08], Bridgeland showed that $S^{[3]}$ is also isomorphic to the moduli space $M_{\sigma}(v)$ of $\sigma$-stable objects in $D^{b}(S)$, where $\sigma$ is a stability condition contained in the distinguished connected component $\operatorname{Stab}^{\dagger}(S)$ (see [Bay18, Section 3] for the definition).
Contractions of moduli spaces of sheaves on K3 surfaces are described in [BM14a]. Their main result is that the wall and chamber decomposition of the movable cone of $M_{\sigma}(v)$ corresponds to the wall and chamber decomposition induced by $v$ on $\operatorname{Stab}^{\dagger}(S)$.
Moreover, the contraction associated to a certain wall $\mathscr{W}\subseteq\operatorname{Stab}^{\dagger}(S)$ is characterized via [BM14a, Theorem 5.7] in terms of the associated hyperbolic lattice $H_{\mathscr{W}}\in H^{*}_{\operatorname{alg}}(S,Z)$ (see [BM14a, Proposition 5.1] for the definition of $H_{\mathscr{W}}$). By definition, the hyperbolic lattice $H_{\mathscr{W}}$ contains $v$, and $\mathscr{W}$ corresponds to the wall $\Theta(H_{\mathscr{W}}^{\perp}\cap v^{\perp})$ of the movable cone of $M_{\sigma}(v)$.
In particular, in our case, contractions of $S^{[3]}$ are given by the two extremal rays ${\bf R}_{\geq 0}L$ and ${\bf R}_{\geq 0}(2L-\delta)$ of $\operatorname{Nef}(S)$. The extremal ray ${\bf R}_{\geq 0}L$ is orthogonal to the isotropic class $w\coloneqq(0,0,-1)$, which satisfies $w\cdot v=1$. The induced contraction is the Hilbert–Chow morphism $S^{[3]}\to S^{(3)}$ ([BM14a, Theorem 5.7(a)]).
@@
$\pi|_{\Delta}\colon\Delta\longrightarrow T,$
where the smooth hyper-Kähler fourfold $T$ is the moduli space of stable sheaves on $S$ with Mukai vector (0,L,4)(0,L,-4).
@@
We use the notation of [BM14a]. Under the isomorphism $\Theta$, the class $2L-\delta$ is the image of the vector w=(1,2L,2)w=(1,-2L,2). In particular, the hyperbolic lattice associated to the contraction induced by $2L-\delta$ is $H=\langle v,(1,-L,2)\rangle=w^{\perp}\subseteq H^{2}_{\operatorname{alg}}(S,{\bf Z})$. It is a lattice with intersection form $\begin{pmatrix}4&0\\ 0&-2\end{pmatrix}$, and contains the spherical class s=(1,L,2)s=(1,-L,2).
@@
Note that, by [Bri08, Lemma 6.2], every $(\alpha,\beta)\in{\bf R}_{>0}\times{\bf R}\smallsetminus\{(1,-1\}$ that satisfies Equation (18) defines a stability condition, hence the wall $\mathscr{W}$ is equal to
@@
We analyze the wall $\mathscr{W}$ using [BM14a, Theorem 5.7]. We first remark that $H$ contains no isotropic classes: indeed a vector $u=xv+ys\in H$, with $x,y\in{\bf Z}$, has square $4x^{2}-2y^{2}$ which is never 0.
Moreover, the wall $\mathscr{W}$ is not totally semistable. Since $\mathscr{W}$ is nonisotropic, to check this, we only need to show that there are no effective spherical classes $u$ with $(u,v)<0$. A spherical class $u\in H$ is a vector $u=xv+ys$ with $x,y\in{\bf Z}$ such that $2x^{2}-y^{2}=-1$, and the condition $(u,v)<0$ is equal to $x<0$. By [BM14a, Proposition 5.5], a class $u$ is effective if and only if $\mathop{\rm Re}\nolimits\left(\frac{Z_{\alpha,\beta}(u)}{Z_{\alpha,\beta}(v)}\right)>0$ for one (hence for all) $(\alpha,\beta)\in\mathscr{W}$. In particular, we can choose $\beta=-2$ and $\alpha=\sqrt{2}$.
Using [Bay18, Equation (3)], we compute $Z_{\alpha,\beta}(v)=-2\beta(\beta+2+i\alpha)$ and $Z_{\alpha,\beta}(s)=-2(\beta+1)(\beta+2+i\alpha)$ for $(\alpha,\beta)\in\mathscr{W}$. We obtain
@@
Note that for each pair $(x,y)$ with $x<0$ that satisfies $2x^{2}-y^{2}=-1$, we have $x+\frac{y}{2}<0$, thus there are no effective spherical classes $u$ with $(u,v)<0$.
Therefore the wall $\mathscr{W}$ is nonisotropic not totally semistable, and the lattice $H$ contains the spherical class $s$, which is effective on $\mathscr{W}$ since $\frac{Z_{\sqrt{2},-2}(s)}{Z_{\sqrt{2},-2}(v)}=1/2$. By [BM14a, Theorem 5.7(a)], the associated contraction $\pi$ is divisorial of Brill–Noether type. In particular, by [BM14a, Lemma 7.1], $\pi$ contracts a divisor $\Delta$ of class $\Theta(s)=L-\delta$. Moreover, since $s$ is the Mukai vector of the line bundle $\mathscr{O}_{S}(-L)$, the divisor $\Delta$ is given by the condition $\operatorname{Hom}(\mathscr{O}_{S}(-L),\bullet)\neq 0$. In particular, this corresponds to length-3 subschemes $Z\subseteq S$ that are contained in some curve $C\in|L|$.
More precisely, following the proof of [BM14a, Lemma 7.4], the image $T\coloneqq\pi(\Delta)$ is isomorphic to the moduli space $M_{\sigma_{0}}(a)$, where $a\coloneqq v-s=(0,L,-4)$ and $\sigma_{0}=\sigma_{\alpha_{0},\beta_{0}}$ is a stability condition lying on the wall $\mathscr{W}$. By [BM14b, Example 9.7], we have equality $M_{\sigma_{0}}(a)=M_{S}(a)$ (by choosing $A=1$, $B=\beta_{0}$), hence $T$ is a smooth hyper-Kähler fourfold that parametrizes sheaves of the form $i_{C,*}(\xi)$, where $i_{C}\colon C\hookrightarrow S$ is the inclusion of a curve $C\in|L|$ in $S$ and $\xi$ is a torsion-free rank $1$ sheaf on $C$ of degree 3-3.
Finally, the fiber of $\pi$ over a point $A]\in M_{S}(a)$ is isomorphic to ${\bf P}(\operatorname{Ext}^{1}(\mathscr{O}_{S}(-L),A))$. Since
@@
If $C\in|L|$ is smooth (of genus 2), a torsion-free rank $1$ sheaf on $C$ of degree 3-3 is of the form $\mathscr{O}_{C}(-Z)$, with $Z\in C^{(3)}$. If $A\coloneqq i_{C,*}(\mathscr{O}_{C}(-Z))$, the preimage $\pi^{-1}([A])$ is equal to $|\mathscr{O}_{C}(Z)|$ (they have the same dimension by Riemann–Roch, and any degree $3$ effective divisor on $C$ equivalent to $Z$ defines a point in $S^{[3]}$ that is sent to $[A]$ via $\pi$). ∎
We keep the notation of the proof. Let $\sigma_{0}$ be a stability condition on the wall $\mathscr{W}$ defined in Equation (19). The moduli space $M_{\sigma_{0}}(v)$ has been constructed as a good moduli space in the sense of Alper (see [Taj23, Theorem 3.3]). By [BM14b, Proposition 8.1], there exists a morphism